Part A A transformer is intended to decrease the value of the alternating voltage from 500...

Part A

A transformer is intended to decrease the value of the alternating voltage from 500 volts to 25 volts. The primary coil contains 200 turns. Find the necessary number of turns N2 in the secondary coil.

Part B

A transformer is intended to decrease the value of the alternating current from 500 amperes to 25 amperes. The primary coil contains 200 turns. Find the necessary number of turns N2 in the secondary coil

Part c:

In a transformer, the primary coil contains 400 turns, and the secondary coil contains 80 turns. If the primary current is 2.5 amperes, what is the secondary current I2?

Express your answer numerically in amperes.
Answer #1

a)
use
N2/N1= V2/V1
N2/200= 25/500
N2=10 turns

b)
use
N2/N1= I1/I2
N2/200= 500/25
N2=4000 turns
C)
use
N2/N1= I1/I2
80/400= 2.5/I2
I2= 12.5 A
Answer #3

apply N1/N2 = V1V2 = I2/I1

A. N2 = V2N1/V1

N2 = 25*200/500

N2 =10 turns

---------------------------------------

B. N2 = 200 * 500/25

N2 =4000 turns

--------------------------------------

C. I2 = 2.5 * 400/80

I2 = 12.5 Amps